OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form. PLEASE SHOW STEP BY STEP

OpenStudy (anonymous):

First find the slope using the 2 point can you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

zero slope or unidentified slop

OpenStudy (anonymous):

\[slope = \frac{ 4-(-4) }{ 1-(-1) }\]

OpenStudy (anonymous):

that's all

OpenStudy (anonymous):

Nope you then need to use this slope and one of the points to find the equation. plug in this formula y-y1=m(x-x1) (x1,y1) is taken from one of the points

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

i dont get my head is about to explode they gave me this question out of no where

OpenStudy (anonymous):

can u show me step by step

OpenStudy (anonymous):

So what do you have so far? m (slope)=4, choose the point (1,4), x1=1, y1=4 you should have y-4=4(x-1)

OpenStudy (anonymous):

can you please just put all the steps on one reply like: 1. 2. 3. 4.

OpenStudy (anonymous):

Step 1: Find the slope(m) = (4-(-4))/(1-(-1)) = 8/2=4

OpenStudy (anonymous):

ok step 2

OpenStudy (anonymous):

Step 2: Use point slope form to get equation: y-y1=m(x-x1), m=4, point (1,4) where x1=1, y1=4 In the point-slope equation: y-1=4(x-4) Do u understand these step so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Step 3: Slope-intercept form is y=mx+b So we have y-1=4(x-4) Solve for y (that is put y by itself on one side).

OpenStudy (anonymous):

If you did that correctly you should get: y=4x-15

OpenStudy (anonymous):

ohhh thanks man

OpenStudy (anonymous):

ill give u a medal

OpenStudy (anonymous):

Sure no problem.